What is the general formula for the pattern 1;9;25;49
Found 2 solutions by Fombitz, amfagge92:
Answer by Fombitz(32388)

Looks like squares of the first 4 odd integers : 1,3,5,7.

Answer by amfagge92(93)

(2n-a)^2 is the general formula,
Where "n" is the term and "a" 1st term,
check:
2*1-1=1Square=1
2*2-1=3 " " " =9
2*3-1=5 " " "=25
2*4-1=7 " " "=49
Thus,1,9,25,49.